OpenStudy (anonymous):

how much is 2x+3y-10 if a) 2x+3y-4=25 b)4x+6y=18?

Parth (parthkohli):

Hmm. This is easy (very) when you think a little about it.

Parth (parthkohli):

\[2x + 3y - 10 \implies 2x + 3y - 4 - 6\]Do you agree with the above?

Parth (parthkohli):

@sarmandi ?

OpenStudy (anonymous):

ah i dont really understand :(

Parth (parthkohli):

Never mind, just tell me if you agree with that one.

OpenStudy (anonymous):

how did u get -6 ?

Parth (parthkohli):

Here's how:\[2x + 3y \color{red}{- 10}\]\[\implies 2x + 3y \color{red}{-4 - 6}\]

OpenStudy (anonymous):

aha okey i got it :D ty :D

Parth (parthkohli):

And did you get the second?

OpenStudy (anonymous):

no i will try do it :D

Parth (parthkohli):

Good job! :)

OpenStudy (anonymous):

ty man ;)

Parth (parthkohli):

Right, let's do it! What do you get when you multiply 2 to (2x + 3y - 10)?

OpenStudy (anonymous):

4x+6x-10 or`?

OpenStudy (anonymous):

6y*

Parth (parthkohli):

Actually, 4x + 6y - 20... not 10 :)

Parth (parthkohli):

OK, let's leave the 4x + 6y - 20 for a while. Returning back to the equation, what do you get by subtracting 20 from both sides in the following?\[4x + 6y = 18\]

OpenStudy (anonymous):

18-19=-1 ??

Parth (parthkohli):

I repeat: subtract 20 FROM BOTH SIDES in the equation — \(4x + 6y = 18\)

OpenStudy (anonymous):

i dont understand what u mean :/

Parth (parthkohli):

\[4x +6y =18 \implies 4x + 6y - 20 = 18 - 20 \]Do you agree?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Okay, simplifying the right-hand-side,\[4x + 6y - 20 = -2\]Now we know that \(2(2x + 3y - 10) = 4x + 6y - 20\) right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So,\[2(2x + 3y - 10) = 2(-1)\]Agree?

OpenStudy (anonymous):

how u got -1?

Parth (parthkohli):

Well, isn't \(-2 = 2(-1)\)?

OpenStudy (anonymous):

answer in book is -1

Parth (parthkohli):

Yeah, the book is right ;)

Parth (parthkohli):

We have \(2(2x + 3y -10) = 2(-1)\). Cancel 2 from both sides!

OpenStudy (anonymous):

ty its right

Parth (parthkohli):

=)

OpenStudy (phi):

how much is 2x+3y-10 if b)4x+6y=18 I think I would factor 2 from the left side in (b) to get 2(2x+3y)= 18 (hoping it will match the question. It looks like it does!) divide both sides of the equation by 2 2x+3y= 9 that looks helpful. replace 2x+3y in the original question with 9: 2x+3y-10 becomes 9-10 or simply -1